OpenStudy (anonymous):

The equation 3^(x-1) + 5^(x-1) = 34 has hw many solutions?

OpenStudy (anonymous):

how about u solve it bro

OpenStudy (anonymous):

No idea lol,

OpenStudy (anonymous):

use wolfram

OpenStudy (anonymous):

No...wolfram

OpenStudy (anonymous):

if \(x>3\) then \( 3^{x-1} + 5^{x-1}>9+25=34\) if \(x<3\) then \( 3^{x-1} + 5^{x-1}<9+25=34\) only answer is \(x=3)

OpenStudy (anonymous):

hm....the answer should be no solution!

OpenStudy (anonymous):

but \(x=3\) satisfies the equation

OpenStudy (anonymous):

@mukushla Can u expl ur method...

OpenStudy (anonymous):

\(3^x\) and \(5^x\) both are increasing functions...

OpenStudy (anonymous):

any other method such taking ex x-2 =t like that

OpenStudy (experimentx):

yep ... that's the only way i know ... analytic approach

OpenStudy (anonymous):

.the answer should be no solution!

OpenStudy (anonymous):

impossible ....... let \(x=3\) then \(3^{3-1} + 5^{3-1}=3^2+5^2=9+25=34\)

OpenStudy (anonymous):

May be my book is wrong! thxx

hero (hero):

lol

OpenStudy (anonymous):

also...lol..:D